I am having trouble finding the solution set on problems like this:
x+y+z=2
3x-y-4z=31
-x-y-5z=-14

Answer by Alan3354(69443)

I am having trouble finding the solution set on problems like this:
x+y+z=2
3x-y-4z=31
-x-y-5z=-14
----------------
Add eqn 1 to eqn 3.
-----
etc.
--------
If you need more help, email via the Thank You note.
Don't bother emailing if you don't do what I said above.